### v4

@@ 5 Isomorphism for Groups of Almost All Orders
Dietrich & Wilson [DW22] previously established that there exists a dense set $\Upsilon\subseteq\mathbb{N}$ such that if $n\in\Upsilon$ and $G_{1},G_{2}$ are magmas of order $n$ given by their multiplication tables, we can (i) decide if $G_{1},G_{2}$ are groups, and (ii) if so, decide whether $G_{1}\cong G_{2}$ in time $O(n^{2}\log^{2}n)$, which is quasi-linear time relative to the size of the multiplication table.
@@
Let $n\in\Upsilon$, and let $G_{1},G_{2}$ be groups of order $n$. We can decide isomorphism between $G_{1}$ and $G_{2}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
@@
$\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, which is contained within $\mathsf{L}\cap\mathsf{FOLL}\subsetneq\textsf{P}$ and cannot compute Parity. While it is known that Group Isomorphism belongs to complexity classes such as $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\textsf{FOLL}$ [CTW13] and $\mathsf{quasiAC^{0}}$ (Theorem 6.1) that cannot compute Parity, membership within P—let alone a subclass of P that cannot compute Parity—is a longstanding open problem.
Proof of Theorem 5.1.
Dietrich & Wilson showed [DW22, Theorem 2.5] that if $G$ is a group of order $n\in\Upsilon$, then $G=H\ltimes B$, where:
@@
Fix a group $G$ of order $n$. We will first discuss how to decompose $G=H\ltimes B$, as prescribed by [DW22, Theorem 2.5]. Without loss of generality, suppose that $p_{1},\ldots,p_{\ell}$ ($\ell\leq k$) are the unique primes where $p_{i}$ ($i\in[\ell]$) divides $n$ only once and $p_{i}>\log\log n$. Now as each $p_{i}$ can be represented as a string of length $\leq\lceil\log(n)\rceil$, we may in $\mathsf{AC}^{0}$ compute $\overline{p}:=p_{1}\cdots p_{\ell}$ (Thm. 2.2). Using Lem. 3.1, we may in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ identify an element $g\in G$ of order $\overline{p}$.
@@
Now in $\mathsf{AC^{0}}$, we may write down the multiplication table for $H_{j}\cong G_{j}/B_{j}$. As $|H_{j}|\leq(\log n)^{\operatorname{poly}\log\log n}$, there are $\operatorname{poly}(n)$ possible generating sequences for $H_{j}$ of length at most $\log\left|\mathinner{H_{j}}\right|$. By [BS84] it actually suffices to consider cube generating sequences for $H_{j}$. Now given cube generating sequences $\overline{x_{j}}$ for $H_{j}$, we may by the proof of Theorem 6.1 decide whether $\overline{x_{1}}\mapsto\overline{x_{2}}$ extends to an isomorphism of $H_{1}$ and $H_{2}$ in $\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. As there are only $\operatorname{poly}(n)$ such generating sequences to consider, we may decide whether $H_{1}\cong H_{2}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Suppose $H_{1}\cong H_{2}$, $B_{1}\cong B_{2}$, and $\text{gcd}(|B_{j}|,|H_{j}|)=1$ for $j=1,2$. We have by the Schur–Zassenhaus Theorem that $G_{j}=H_{j}\ltimes_{\theta_{j}}B_{j}$ ($j=1,2$). By Taunt’s Lemma [Tau55], it remains to test whether the actions $\theta_{1}$ and $\theta_{2}$ are equivalent in the following sense. We first note that, as $B_{j}$ is Abelian, for any two elements $h_{1},h_{2}$ of $G_{j}$ belonging to the same coset of $G_{j}/B_{j}$ and any element $b\in B_{j}$, that $h_{1}bh_{1}^{-1}=h_{2}bh_{2}^{-1}$. Thus, given an isomorphism $\alpha:H_{1}\cong H_{2}$ (where $\alpha(\overline{h_{1}})=\overline{h_{2}}$ is represented using arbitrary coset representatives of $\overline{h_{1}},\overline{h_{2}}$), we may by Lemma 3.1, search in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ for generators $b_{1}\in B_{1},b_{2}\in B_{2}$ and write down the isomorphism $\beta:B_{1}\cong B_{2}$ induced by the map $b_{1}\mapsto b_{2}$. We may now check in $\mathsf{AC^{0}}$ whether for all $\overline{h_{1}}\in H_{1}$ and all $b\in B_{j}$ whether
$\beta(\overline{h_{1}}b\overline{h_{1}^{-1}})=\alpha(\overline{h_{1}})\beta(b)\alpha(\overline{h_{1}^{-1}}).$
Here, we abuse notation, where $\overline{h_{1}}b\overline{h_{1}^{-1}}$ is evaluated using the coset representatives $h_{1}\in\overline{h_{1}}$ and $h_{1}^{-1}\in\overline{h_{1}^{-1}}$. The result now follows. ∎
@@ 6 Quasigroup Isomorphism
Quasigroup Isomorphism belongs to $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. In particular, it can be solved by $\mathsf{quasiAC^{0}}$ circuits of depth 4.
Note that we have $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))\subseteq\mathsf{quasiAC^{0}}\cap\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$.
This statement is inspired by [DEP+22] where a similar problem is shown to be in the third level of the polynomial-time hierarchy using the same approach. Our proof follows closely the algorithm for [CTW13, Theorem 3.4].
Proof of Thm. 6.1.
Let $G$ and $H$ be quasigroups given as their multiplication tables. We assume that the elements of the quasigroups are indexed by integers $1,\dots,\left|\mathinner{G}\right|$. If $\left|\mathinner{G}\right|\neq\left|\mathinner{H}\right|$ (this can be tested in $\mathsf{DTIME}(\log n)$ by a standard binary search), we know that $G$ and $H$ are not isomorphic. Otherwise, let us write $n=\left|\mathinner{G}\right|$ and $k=\left\lceil\mathinner{2\log n}\right\rceil+1$ (while [CTW13, Theorem 3.3] only states that some $k\in O(\log n)$ suffices, the corresponding proof states that for $k=\left\lceil\mathinner{2\log n}\right\rceil+1$, there is actually a cube generating set).
The basic idea is to guess cube generating sequences $(g_{0},\dots,g_{k})$ and $(h_{0},\dots,h_{k})$ for $G$ and $H$ and verify that the map $g_{i}\mapsto h_{i}$ induces an isomorphism between $G$ and $H$. Hence, we start by guessing cube generating sequences $(g_{0},\dots,g_{k})$ and $(h_{0},\dots,h_{k})$ with respect to the parenthesization $P$ where the elements are evaluated left-to-right (so $P(g_{1}g_{2}g_{3})=(g_{1}g_{2})g_{3}$), with $g_{i}\in G$, $h_{i}\in H$. This amounts to guessing $k\cdot\log(n)\in O(\log^{2}n)$ many bits (thus, $\exists^{\log^{2}n}$). Now, we need to verify two points in $\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$:
@@
Let us describe the first point for $G$ (for $H$ this follows exactly the same way): we universally verify for every element $g\in G$ (which can be encoded using $O(\log n)$ bits, hence, $\forall^{\log n}$) that we can existentially guess a sequence $(e_{1},\dots,e_{k})\in\{0,1\}^{k}$ (i.e. $\exists^{\log n}$) such that $g=P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$. We can compute this product in $\mathsf{DTISP}(\log^{2+o(1)}n,\log n)$ by multiplying from left to right. [^16] Here, we could impose the additional requirement that the length of each row/column of the multiplication table is padded up to a power of two in order to get a bound of $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\log^{2}n,\log n)$) Each multiplication can be done in time $O(\log^{1+o(1)}n)$ because we simply need to compute $i+j\cdot n$ for two addresses $i,j$ of quasigroup elements, write the result on the index tape and then read the corresponding product of group elements from the multiplication table. Moreover, note that for this procedure we only need to store one intermediate result on the working tape at any time. Thus, computing the product $P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$ can be done in time $O(\log^{2+o(1)}n)$ and space $O(\log n)$ and it can be checked whether the result is $g$.
To check the second point, by [CTW13], we need to verify universally that for all $(c_{1},\dots,c_{k})$, $(d_{1},\dots,d_{k})$, $(e_{1},\dots,e_{k})\in\{0,1\}^{k}$ (hence, $\forall^{\log n}$) whether
$P(g_{0}g_{1}^{c_{1}}\cdots g_{k}^{c_{k}})\cdot P(g_{0}g_{1}^{d_{1}}\cdots g_{k}^{d_{k}})=P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}}).$
Again the products can be computed in time $O(\operatorname{poly}\log n)$ and space $O(\log n)$ as outlined above. Now, it remains to observe that we can combine the two $\forall^{\log n}$ blocks into one $\forall^{\log n}$ block because the check whether $g_{i}\mapsto h_{i}$ induces an isomorphism does not depend on the existentially guessed bits in the first check.
The depth 4 circuit bound then follows from the DNF for $\mathsf{DTIME}(\text{polylog}(n))$ (Fact 2.4): the quantifier blocks $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}$ turn into $\bigvee^{\log^{2}n}\bigwedge^{\log n}\bigvee^{\log n}$, and then the disjunction at the top of the DNF for $\mathsf{DTIME}(\text{polylog}(n))$ can be merged into the final $\bigvee$ from the quantifier blocks. ∎
@@
    Latin Square Isotopy. In particular, this problem belongs to $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which yields depth $4$ $\mathsf{quasiAC^{0}}$ circuits.
@@
    Isomorphism testing of two Steiner triple systems. In particular, this problem belongs to
    $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which yields depth $4$ $\mathsf{quasiAC^{0}}$ circuits.
@@
For Latin Square Isotopy and isomorphism testing of Steiner triple systems, a careful analysis yields depth-$4$ $\mathsf{quasiAC^{0}}$ circuits. In contrast, the reductions from [Lev23] for isomorphism testing of Latin square graphs, Steiner $(t,t+1)$-designs, and pseudo-STS graphs all use circuits of depth at least $4$. Obtaining reductions of depth less than $4$ will likely require new techniques.
Proof.
@@
    We note that the reductions outlined in [Mil78, Theorem 2] and [Lev23, Remark 1.6] in fact allow us to determine whether two quasigroup are isotopic, and not just main class isomorphic. Thus, we have an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction from Latin Square Isotopy to Quasigroup Isomorphism. This suffices to yield the bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
    We now turn to establishing the stronger $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ bound. We carefully analyze the $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ algorithm from [Lev23, Section 3] using cube generating sequences. Our analysis adapts the proof technique of Theorem 6.1. Let $Q_{1},Q_{2}$ be quasigroups, and let $k\in O(\log n)$. We begin by guessing three cube generating sequences $\bar{a}=(a_{0},a_{1},\ldots,a_{k}),\bar{b}=(b_{0},b_{1},\ldots,b_{k}),\bar{c}=(c_{0},c_{1},\ldots,c_{k})$ for $Q_{1}$, and $\bar{a^{\prime}}=(a_{0}^{\prime},a_{1}^{\prime},\ldots,a_{k}^{\prime}),\bar{b^{\prime}}=(b_{0}^{\prime},b_{1}^{\prime},\ldots,b_{k}^{\prime}),\bar{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime},\ldots,c_{k}^{\prime})$ for $Q_{2}$. This requires $O(\log^{2}n)$ non-deterministic bits ($\exists^{\log^{2}n}$). From the proof of Theorem 6.1, we may check in $\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ whether each of $\bar{a},\bar{b},\bar{c}$ generates $Q_{1}$, and whether $\bar{a^{\prime}},\bar{b^{\prime}},\bar{c^{\prime}}$ generates $Q_{2}$.
@@
    By Fact 2.4, an $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ machine can be simulated by a depth-$2$ $\mathsf{quasiAC^{0}}$ circuit, taking care of the final $\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. The additional quantifiers $\exists^{\log^{2}n}\forall^{\log n}$ then yield depth-$4$ $\mathsf{quasiAC^{0}}$ circuits.
@@
    Given two Latin square graphs $G_{1},G_{2}$, we may recover corresponding Latin squares $L_{1},L_{2}$ in $\mathsf{AC^{0}}$ (cf. [Lev23, Lemma 3.9]). Now $G_{1}\cong G_{2}$ if and only if $L_{1}$ and $L_{2}$ are main class isomorphic [Mil78, Lemma 3]. We may decide whether $L_{1},L_{2}$ are main class isomorphic using an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction to Quasigroup Isomorphism (cf. [Lev23, Remark 1.6]). By Fact 2.7, $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is closed under $\mathsf{AC^{0}}$-computable dtt reductions, thus yielding the bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
@@
    This immediately follows from the fact that isomorphism testing of Steiner $(t,t+1)$-designs is $\exists^{\log^{2}n}\mathsf{AC^{0}}$-reducible to isomorphism testing of Steiner triple systems [BW13] (cf., [Lev23, Corollary 4.11]).

### v5

@@ 5 Isomorphism for Groups of Almost All Orders
Dietrich & Wilson [22] previously established that there exists a dense set $\Upsilon\subseteq\mathbb{N}$ such that if $n\in\Upsilon$ and $G_{1},G_{2}$ are magmas of order $n$ given by their multiplication tables, we can (i) decide if $G_{1},G_{2}$ are groups, and (ii) if so, decide whether $G_{1}\cong G_{2}$ in time $O(n^{2}\log^{2}n)$, which is quasi-linear time relative to the size of the multiplication table.
@@
Let $\Upsilon\subseteq\mathbb{N}$ be the dense set considered by Dietrich and Wilson [22]. Let $n\in\Upsilon$, and let $G_{1},G_{2}$ be groups of order $n$. We can decide isomorphism between $G_{1}$ and $G_{2}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
@@
$\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$, which is contained within $\mathsf{L}\cap\mathsf{FOLL}\subsetneq\mathsf{P}$ and cannot compute Parity. While it is known that Group Isomorphism belongs to complexity classes such as $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ [17] and $\mathsf{quasiAC^{0}}$ (Theorem 6.1) that cannot compute Parity, membership within $\mathsf{P}$—let alone a subclass of $\mathsf{P}$ that cannot compute Parity—is a longstanding open problem.
Proof 5.3 (Proof of Theorem 5.1).
Dietrich & Wilson showed [22, Theorem 2.5] that if $G$ is a group of order $n\in\Upsilon$, then $G=H\ltimes B$, where $\gcd(|B|,|H|)=1$ and:
@@
Fix a group $G$ of order $n$. We will first discuss how to decompose $G=H\ltimes B$, as prescribed by [22, Theorem 2.5]. Without loss of generality, suppose that $p_{1},\ldots,p_{\ell}$ ($\ell\leq k$) are the unique primes where $p_{i}$ ($i\in[\ell]$) divides $n$ only once and $p_{i}>\log\log n$. Now as each $p_{i}$ can be represented as a string of length $\leq\lceil\log(n)\rceil+1$, we may in $\mathsf{AC}^{0}$ compute $\overline{p}:=p_{1}\cdots p_{\ell}$ (Thm. 2.2). Using Lem. 3.1, we may in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ identify an element $g\in G$ of order $\overline{p}$.
@@
Now in $\mathsf{AC^{0}}$, we may write down the multiplication table for $H_{j}\cong G_{j}/B_{j}$. As $|H_{j}|\leq(\log n)^{\operatorname{poly}\log\log n}$, there are $\operatorname{poly}(n)$ possible generating sequences for $H_{j}$ of length at most $\log\left|\mathinner{H_{j}}\right|$. By [7] it actually suffices to consider cube generating sequences for $H_{j}$. Now given cube generating sequences $\overline{x_{j}}$ for $H_{j}$, we may by the proof of Theorem 6.1 decide whether $\overline{x_{1}}\mapsto\overline{x_{2}}$ extends to an isomorphism of $H_{1}$ and $H_{2}$ in $\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))\subseteq\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$. As there are only $\operatorname{poly}(n)$ such generating sequences to consider, we may decide whether $H_{1}\cong H_{2}$ in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
Suppose $H_{1}\cong H_{2}$, $B_{1}\cong B_{2}$, and $\text{gcd}(|B_{j}|,|H_{j}|)=1$ for $j=1,2$. We have by the Schur–Zassenhaus Theorem that $G_{j}=H_{j}\ltimes_{\theta_{j}}B_{j}$ ($j=1,2$) for some action $\theta_{j}\colon H_{j}\to\text{Aut}(B_{j})$. By Taunt’s Lemma [54], it remains to test whether the actions $\theta_{1}$ and $\theta_{2}$ are equivalent in the following sense: do there exist isomorphisms $\alpha\colon H_{1}\cong H_{2}$ and $\beta\colon B_{1}\cong B_{2}$ such that
$\beta(hbh^{-1})=\alpha(h)\beta(b)\alpha(h^{-1})\qquad\forall h\in H_{1},b\in B_{1}?$
Note that, as $B_{j}$ is Abelian, for any two elements $h_{1},h_{2}$ of $G_{j}$ belonging to the same coset of $B_{j}$ and any element $b\in B_{j}$, that $h_{1}bh_{1}^{-1}=h_{2}bh_{2}^{-1}$, so the above conjugation action is well-defined, independent of the choice of isomorphic copy of $H_{j}$ in $G_{j}$. Next, since the $B_{j}$ are cyclic, if $b_{1}$ generates $B_{1}$, then the above condition is satisfied iff for all $h\in H_{1}$ we have
$\beta(hb_{1}h^{-1})=\alpha(h)\beta(b_{1})\alpha(h^{-1})\qquad\forall h\in H_{1}.$
As above, in $\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ we can find a generator $b_{1}\in B_{1}$ and a cube generating sequence $h_{1},\dotsc,h_{k}\in H_{1}$. In parallel, for all $\operatorname{poly}(n)$ generators $b_{2}\in B_{2}$ and all $\operatorname{poly}(n)$ cube generating sequences $h^{\prime}_{1},\dotsc,h^{\prime}_{k}\in H_{2}$, we then test the above condition on the isomorphisms specified by $\beta(b_{1})=b_{2}$ and $\alpha(h_{i})=h^{\prime}_{i}$ for $i=1,\dotsc,k$. That is, for each choice of $b_{2}$ we write down the isomorphism $B_{1}\to B_{2}$ defined by $\beta(b_{1})=b_{2}$, and for each choice of cube generating sequence $h^{\prime}_{1},\dotsc,h^{\prime}_{k}\in H_{2}$, we check that
$\beta(h_{i}b_{1}h_{i}^{-1})=h^{\prime}_{i}b_{2}(h^{\prime}_{i})^{-1}\qquad\forall i=1,\dotsc,k.$
Checking the preceding condition only involves a constant number of group multiplications, which can thus be done in $\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$.
@@ 6 Quasigroup Isomorphism
Quasigroup Isomorphism belongs to $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. In particular, it can be solved by $\mathsf{quasiAC^{0}}$ circuits of depth 4 and size $n^{O(\log n)}$.
Note that we have
$\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ $\subseteq\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$
$\subseteq\mathsf{quasiAC^{0}}\cap\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}.$
This statement is inspired by [21] where a similar problem is shown to be in the third level of the polynomial-time hierarchy using the same approach. Our proof follows closely the algorithm for [17, Theorem 3.4].
Proof 6.2 (Proof of Theorem 6.1).
Let $G$ and $H$ be quasigroups given as their multiplication tables. We assume that the elements of the quasigroups are indexed by integers $1,\dots,\left|\mathinner{G}\right|$. If $\left|\mathinner{G}\right|\neq\left|\mathinner{H}\right|$ (this can be tested in $\mathsf{DTIME}(\log n)$ by a standard binary search), we know that $G$ and $H$ are not isomorphic. Otherwise, let us write $n=\left|\mathinner{G}\right|$ and $k=\left\lceil\mathinner{2\log n}\right\rceil+1$ (while [17, Theorem 3.3] only states that some $k\in O(\log n)$ suffices, the corresponding proof states that for $k=\left\lceil\mathinner{2\log n}\right\rceil+1$, there is actually a cube generating set).
The basic idea is to guess cube generating sequences $(g_{0},\dots,g_{k})$ and $(h_{0},\dots,h_{k})$ for $G$ and $H$ and verify that the map $g_{i}\mapsto h_{i}$ induces an isomorphism between $G$ and $H$. Hence, we start by guessing cube generating sequences $(g_{0},\dots,g_{k})$ and $(h_{0},\dots,h_{k})$ with respect to the parenthesization $P$ where the elements are evaluated left-to-right (so $P(g_{1}g_{2}g_{3})=(g_{1}g_{2})g_{3}$), with $g_{i}\in G$, $h_{i}\in H$. This amounts to guessing $2k\cdot\log(n)\in O(\log^{2}n)$ many bits (thus, $\exists^{\log^{2}n}$). Now, we need to verify two points in $\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$:
@@
Let us describe the first point for $G$ (for $H$ this follows exactly the same way): we universally verify for every element $g\in G$ (which can be encoded using $O(\log n)$ bits, hence, $\forall^{\log n}$) that we can existentially guess a sequence $(e_{1},\dots,e_{k})\in\{0,1\}^{k}$ (i.e. $\exists^{\log n}$) such that $g=P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$. We can compute this product in $\mathsf{DTISP}(\log^{2+o(1)}n,\log n)$ by multiplying from left to right: [^16] Here, we could impose the additional requirement that the length of each row/column of the multiplication table is padded up to a power of two in order to get a bound of $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\log^{2}n,\log n)$). Each multiplication can be done in time $O(\log^{1+o(1)}n)$ because we simply need to compute $i+j\cdot n$ for two addresses $i,j$ of quasigroup elements, write the result on the index tape and then read the corresponding product of group elements from the multiplication table. Moreover, note that for this procedure we only need to store one intermediate result on the working tape at any time , and one quasigroup multiplication only queries $O(\log n)$ bits (this point will be important for our size analysis below); the “+o(1)+o(1)” in the exponent is merely to perform the arithmetic. Thus, computing the product $P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$ can be done in time $O(\log^{2+o(1)}n)$ and space $O(\log n)$ and it can be checked whether the result is $g$.
To check the second point, by [17], we need to verify universally that for all $(c_{1},\dots,c_{k})$, $(d_{1},\dots,d_{k})$, $(e_{1},\dots,e_{k})\in\{0,1\}^{k}$ (hence, $\forall^{\log n}$) whether
$P(g_{0}g_{1}^{c_{1}}\cdots g_{k}^{c_{k}})\cdot P(g_{0}g_{1}^{d_{1}}\cdots g_{k}^{d_{k}})=P(g_{0}g_{1}^{e_{1}}\cdots g_{k}^{e_{k}})$
$\iff$ $P(h_{0}h_{1}^{c_{1}}\cdots h_{k}^{c_{k}})\cdot P(h_{0}h_{1}^{d_{1}}\cdots h_{k}^{d_{k}})=P(h_{0}h_{1}^{e_{1}}\cdots h_{k}^{e_{k}})$
These products can be computed in the same asymptotic complexity bounds as the product above, using the same technique. Now, it remains to observe that we can combine the two $\forall^{\log n}$ blocks into one $\forall^{\log n}$ block because the check whether $g_{i}\mapsto h_{i}$ induces an isomorphism does not depend on the existentially guessed bits in the first check.
The depth 4 circuit bound then follows from the DNF for $\mathsf{DTIME}(\text{polylog}(n))$ (Fact 1): the quantifier blocks $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}$ turn into $\bigvee^{\log^{2}n}\bigwedge^{\log n}\bigvee^{\log n}$, and then the disjunction at the top of the DNF for $\mathsf{DTIME}(\text{polylog}(n))$ can be merged into the final $\bigvee$ from the quantifier blocks. The size bound is computed as follows: the $\bigvee^{\log^{2}n}\bigwedge^{\log n}\bigvee^{\log n}$ multiplies the size by $2^{O(\log^{2}n+2\log n)}$. For each of the above two points, we showed it can be decided in $\mathsf{DTIME}((\log n)^{2+o(1)})$ using at most $O(\log^{2}n)$ queries to the input. By the proof of Fact 1, those yield decision trees of depth $O((\log n)^{2})$, giving a DNF of size at most $n\cdot 2^{O(\log n)^{2}}$. Multiplying all these factors together gives size $2^{O(\log^{2}n)}=n^{O(\log n)}$.
@@
    Latin Square Isotopy. In particular, it belongs to
    $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which yields depth 44 $\mathsf{quasiAC^{0}}$ circuits.
@@
    Isomorphism testing of Steiner triple systems. In particular, this problem belongs to
    $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$, which yields depth 44 $\mathsf{quasiAC^{0}}$ circuits.
@@
For Latin Square Isotopy and isomorphism testing of Steiner triple systems, a careful analysis yields depth-$4$ $\mathsf{quasiAC^{0}}$ circuits. In contrast, the reductions from [37] for isomorphism testing of Latin square graphs, Steiner $(t,t+1)$-designs, and pseudo-STS graphs all use circuits of depth at least $4$. Obtaining reductions of depth less than $4$ will likely require new techniques.
Proof 6.4.
@@
    We note that the reductions outlined in [43, Theorem 2] and [37, Remark 1.6] in fact allow us to determine whether two quasigroups are isotopic, and not just main class isomorphic. Thus, we have an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction from Latin Square Isotopy to Quasigroup Isomorphism. This suffices to yield the bound of $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
    We now turn to establishing the stronger $\exists^{\log^{2}n}\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ bound. We carefully analyze the $\exists^{\log^{2}n}\mathsf{L}\cap\exists^{\log^{2}n}\mathsf{FOLL}$ algorithm from [37, Section 3] using cube generating sequences. Our analysis adapts the proof technique of Theorem 6.1. Let $Q_{1},Q_{2}$ be quasigroups, and let $k\in O(\log n)$. We begin by guessing three cube generating sequences $\bar{a}=(a_{0},a_{1},\ldots,a_{k}),\bar{b}=(b_{0},b_{1},\ldots,b_{k}),\bar{c}=(c_{0},c_{1},\ldots,c_{k})$ for $Q_{1}$, and $\bar{a^{\prime}}=(a_{0}^{\prime},a_{1}^{\prime},\ldots,a_{k}^{\prime}),\bar{b^{\prime}}=(b_{0}^{\prime},b_{1}^{\prime},\ldots,b_{k}^{\prime}),\bar{c^{\prime}}=(c_{0}^{\prime},c_{1}^{\prime},\ldots,c_{k}^{\prime})$ for $Q_{2}$. This requires $O(\log^{2}n)$ non-deterministic bits ($\exists^{\log^{2}n}$). From the proof of Theorem 6.1, we may check in $\forall^{\log n}\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ whether each of $\bar{a},\bar{b},\bar{c}$ generates $Q_{1}$, and whether $\bar{a^{\prime}},\bar{b^{\prime}},\bar{c^{\prime}}$ generates $Q_{2}$.
@@
    By Fact 1, an $\mathsf{NTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$ machine can be simulated by a depth-$2$ $\mathsf{quasiAC^{0}}$ circuit, taking care of the final $\exists^{\log n}\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n))$. The additional quantifiers $\exists^{\log^{2}n}\forall^{\log n}$ then yield depth-$4$ $\mathsf{quasiAC^{0}}$ circuits.
@@
    Given two Latin square graphs $G_{1},G_{2}$, we may recover corresponding Latin squares $L_{1},L_{2}$ in $\mathsf{AC^{0}}$ (cf. [37, Lemma 3.9]). Now $G_{1}\cong G_{2}$ if and only if $L_{1}$ and $L_{2}$ are main class isomorphic [43, Lemma 3]. We may decide whether $L_{1},L_{2}$ are main class isomorphic using an $\mathsf{AC^{0}}$-computable disjunctive truth-table reduction to Quasigroup Isomorphism (cf. [37, Remark 1.6]). By Fact 3, $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$ is closed under $\mathsf{AC^{0}}$-computable dtt reductions, thus yielding the bound of
    $\exists^{\log^{2}n}\mathsf{AC^{0}}(\mathsf{DTISP}(\mathrm{polylog}(n),\mathrm{log}(n)))$.
@@
    This immediately follows from the fact that isomorphism testing of Steiner $(t,t+1)$-designs is $\exists^{\log^{2}n}\mathsf{AC^{0}}$-reducible to isomorphism testing of Steiner triple systems [8] (cf., [37, Corollary 4.11]).